### v2

Abstract
The Great Oxidation Event (GOE) was a $200$ Myr transition circa 2.4 billion years ago that converted the Earth’s anoxic atmosphere to one where molecular oxygen (O2) was abundant (volume mixing ratio $10^{-4}$). This significant rise in O2 is thought to have substantially throttled hydrogen (H) escape and the associated water (H2O) loss. Atmospheric estimations from the GOE onward place O2 concentrations ranging between 0.1% to 150% PAL, where PAL is the present atmospheric level of 21% by volume. In this study we use WACCM6, a three-dimensional Earth System Model to simulate Earth’s atmosphere and predict the diffusion-limited escape rate of hydrogen due to varying O2 post-GOE. We find that O2 indirectly acts as a control valve on the amount of hydrogen atoms reaching the homopause in the simulations: less O2 leads to decreased O3 densities that reduce local tropical tropopause temperatures by up to 17 K, which increases H2O freeze-drying and thus reduces the primary source of hydrogen in the considered scenarios. The maximum differences between all simulations in the total H mixing ratio at the homopause and the associated diffusion-limited escape rates are a factor of 3.2 and 4.7, respectively. The prescribed CH4 mixing ratio (0.8 ppmv) sets a minimum diffusion escape rate of $\approx 2\times 10^{10}$ mol H yr-1, effectively a negligible rate when compared to pre-GOE estimates ($\sim 10^{12}-10^{13}$ mol H yr-1). Because the changes in our predicted escape rates are comparatively minor, our numerical predictions support geological evidence that the majority of Earth’s hydrogen escape occurred prior to the GOE. Our work demonstrates that estimations of how the tropical tropopause layer and the associated hydrogen escape rate evolved through Earth’s history requires 3D chemistry-climate models which include a global treatment of water vapour microphysics.
[…]
[Introduction]
A key component of Earth’s atmospheric and biological history is its water (H2O). The relative amount of H2O that was present during the Earth’s formation versus the quantity delivered by volatile-rich bodies (e.g., comets and asteroids; 2000Icar..148..508D; 2013ApJ...767...54I), as well as the H2O lost to space through hydrogen escape (2001Sci...293..839C; 2008Icar..194...42G), or exchanged into the interior (2015Sci...350..795H; 2017RSPTA.37550393K), has had an important impact on how the Earth’s surface and the atmosphere have developed. Beyond Earth, the solar system’s terrestrial planet atmospheres are desiccated and consequently, Venus and Mars are both apparently desolate environments. It’s possible that Venus was never habitable (2021Natur.598..276T; 2023PNAS..12009751W; 2024NatAs.tmp..289C), but Mars was likely habitable early in its history before its water was isolated to the subsurface and the crust (2010NatGe...3..459D; 2018Sci...361..490O; 2021Sci...372...56S; 2021NatAs...5...63L; 2022NatAs...6.1263S; 2024RemS...16..824Z; wright2024liquid). Venus is now too hot for liquid H2O to persist and probably lost much of its natal H2O through atmospheric hydrogen escape (1969JAtS...26.1191I; 2006P&SS...54.1425K; 2019JGRE..124.2015K; 2024AGUFMSM13E..08C; Gu_Venus_Photo_2025).
In contrast, Earth has mostly remained temperate throughout its history, with liquid water persisting on the surface for over 4 billion years (2012RvGeo..50.2006F, and references therein), although there were periods of widespread glaciation at either end of the Proterozoic eon (2.4 – 0.541 Gyr ago; young2013precambrian; 2020PNAS..11713314W). A mostly temperate Earth has been possible despite a fainter Sun in Earth’s past (bahcall2001solar; 2012RvGeo..50.2006F; 2020SSRv..216...90C), in combination with fluctuations in albedo (2020SSRv..216...90C; goldblatt2021earth) and atmospheric composition, in particular the greenhouse gases CO2 (2006PreR..147..148S; 2013ChGeo.362..224S; kanzaki2015estimates) and CH4 (daines2016effect; zhao2018terrestrial; 2019E&PSL.522...48L; fakhraee2019proterozoic). Geochemical proxies suggest a broad range but generally greater past abundance for CO2, whereas only loose constraints have been placed on CH4 (2020SciA....6.1420C, and references therein), with debates regarding whether CH4 during the Proterozoic was higher or lower in abundance than present day still ongoing (see Discussion section).
As all known life requires liquid H2O, the history of Earth’s H2O is inextricably tied to the evolution of Earth’s life. From isotopic evidence and computational modelling, since its formation, the Earth may have lost between 0.13 and 2 times the present ocean volume (2012PNAS..109.4371P; 2018E&PSL.497..149K; 2019GeCoA.244...56Z). Indeed, the early Archean Earth (4 – 2.4 Gyr ago), before 3.2 Gyr ago, may have been fully covered in a global ocean (2021AGUA....200323D). The current hydrogen escape rate is $\approx 4\times 10^{10}$ mol of H $\textrm{yr}^{-1}$, corresponding to roughly 1 metre of global water loss per 1 billion years (2013ChGeo.362...26Z). Given that the Earth’s ocean can be up to 11 km deep (gardner2014so), and its mean depth is $\approx 3.6$ km (charette2010volume), the loss rate is insignificant. This suggests a different past for hydrogen escape than the constant modern rate would imply.
Therefore, an important but uncertain component of Earth’s water loss is determining the amount of hydrogen (H) that has escaped to space. As the lightest element, H escapes atmospheres more readily than other elements and chemical compounds. On Earth, hydrogen atoms exist in many different chemical species in the lower and middle atmosphere (2005ama..book.....B). The four major chemical constituents that contribute hydrogen atoms to the upper atmosphere, and therefore the species we focus on in this work, are H, H2, H2O, and CH4. Atmospheric escape of hydrogen can occur from a planet through several mechanisms, including: Jeans escape, hydrodynamic flow, sputtering, impacts, photochemical escape, and charge exchange escape (2007SSRv..129..245L; 2010ApJ...716.1573J; 2018PhyU...61..217S; 2020JGRA..12527639G). For these mechanisms, with the exception of impacts and hydrodynamic flow, the hydrogen escapes from the exosphere at the top of the atmosphere. In such cases, hydrogen cannot escape faster than it can be delivered to the exosphere through molecular diffusion. This is known as ‘diffusion-limited hydrogen escape’ (1973JAtS...30.1481H; 1976AREPS...4..265H; 1977evat.book.....W; 2003ARA&A..41..429K; 2013ChGeo.362...26Z). The diffusion-limited rate of hydrogen escape is set by two controlling bottlenecks, the upward diffusion of hydrogen atoms (mostly in H2O and CH4) through the tropopause ‘cold trap’, and the upward diffusion of hydrogen atoms through molecular diffusion above the homopause [^1] The homopause is the atmospheric point above which molecular diffusion dominates and below which turbulent diffusion dominates (2001Sci...293..839C; catling2017atmospheric).
The primary hydrogen bottleneck, the cold trap, is where the saturation mixing ratio of H2O and ice is at a minimum, such that it modulates the amount of H2O propagating upwards into the stratosphere (2015ApJ...813L...3K). On Earth, the cold trap is typically described as the tropical tropopause layer (TTL). The TTL is three-dimensional and is defined as the transition region between the Hadley circulation in the troposphere and the start of the deep branch of the Brewer-Dobson circulation (2005ama..book.....B; 2009RvGeo..47.1004F; 2017ACP....17.4337G). Despite being located above the tropics (at 12 – 16 km in altitude, which is 70 – 150 hPa in pressure), which have on average greater surface temperatures than higher latitudes, the TTL is defined by its cold temperatures (2005ama..book.....B; 2009RvGeo..47.1004F). The cause of this is adiabatic cooling as air is transported upwards from the convective troposphere into the stratosphere, and this cold temperature region leads to the removal of water vapor from the atmosphere through condensation (1949QJRMS..75..351B; newell1981stratospheric; 1982GeoRL...9..605D; 2007ACPD....7.8933P). The location, temperatures, and dynamics of the TTL are affected by seasons (kim2012tropical) and climate change (2013NatGe...6..169R; 2017JCli...30.1245L; 2023FrEaS..1177502Z; 2024ACP....24.7405Z). Of the major H-bearing species, only H2O is significantly affected by temperatures in the TTL, whereas H2 and CH4, which don’t condense in Earth’s atmosphere, are unaffected by the cold trap mechanism.
In the stratosphere, hydrogen atoms are interchanged between CH4, H2O and H2, until reaching the homopause (catling2017atmospheric). At altitudes exceeding the homopause, through diffusive separation, the lighter atmospheric constituents with lower molecular masses increase in relative abundance with decreasing pressure (e.g., H, He, O, and N). Note that H2 would also increase but it is rapidly photodissociated into H. Hence, above the homopause, H dominates the total hydrogen budget (catling2017atmospheric).
The atmosphere of the Archean Earth was likely weakly reducing with greater quantities of CH4, CO, and H2 (2001OLEB...31..271K; kharecha2005coupled; 2019GeCoA.262..207K; 2020SciA....6.1420C). Midway through Earth’s history, approximately 2.4 billion years ago, a rise in oxygen known as the Great Oxidation Event (GOE; 2000Sci...289..756F; 2002GeCoA..66.3811H; 2019PNAS..11617207H; 2020PNAS..11713314W; 2021NatGe..14..143G) has been proposed to have significantly reduced hydrogen escape from $\sim 10^{12}-10^{13}$ mol H yr-1 to $\sim 10^{10}$ mol H yr-1 (2001Sci...293..839C; 2013ChGeo.362...26Z). This is because an oxygenated atmosphere destroys reducing species (e.g., H2 and CH4), resulting in lower amounts of such gases which contribute hydrogen atoms to the upper atmosphere.
In conjunction, atmospheric escape of hydrogen may have played a significant role in shaping Earth’s atmosphere through time. It is thought to at least be partially responsible for oxidising the Earth’s surface: when H2O is photodissociated above the troposphere, the liberated hydrogen (H) can escape to space, and the oxygen (O) left behind oxidises the Earth (2001Sci...293..839C; claire2006biogeochemical; 2013ChGeo.362...26Z). Alongside photosynthetic production of O2 from cyanobacteria, this process is hypothesized to have contributed to the occurrence of the GOE because the process of hydrogen escape and subsequent oxidation is irreversible, leading to a reduced sink of O2 over time (2001Sci...293..839C; claire2006biogeochemical; 2013ChGeo.362...26Z). Nevertheless, considerable uncertainty persists regarding the precise temporal evolution of Earth’s atmosphere (kasting2025evolution), with hydrogen escape a key aspect of the holistic picture.
There is no direct way to determine the amount of hydrogen that has escaped in Earth’s history. As mentioned, several measurements can be used to infer this quantity, and the chronology of one of the heaviest gases, Xe, may be informative. Xe is more easily ionised compared to the other noble gases which do not exhibit the same fractionation (the so-called Xenon paradox; 1977Sci...198..453A; 2013NatCh...5...61Z). It was proposed that Xe+ could escape Earth’s atmosphere when propelled by H+ ions, and this has been shown to be a feasible physical process (2019GeCoA.244...56Z). Accordingly, the timing of Earth’s Xe fractionation (see figure 3 in 2019GeCoA.244...56Z) may trace high levels of atmospheric hydrogen escape during the Archean eon. There has been relatively small changes in Xe fractionation since the GOE (2018GeCoA.232...82A), indicating the vast majority of historical hydrogen escape occurred prior to this time (i.e., before $\sim 2.4$ Gyr ago).
Whilst hydrogen escape rates may have been much lower in the last $\sim 2.4$ Gyr, to our knowledge, no work has quantitatively estimated escape rates since the GOE occurred using a 3D model. In this work, we build upon a previous study (2022RSOS....911165C) which simulated a range of O2 concentrations that may have been present on the Earth for the last 2.4 Gyr and found that lower O3 columns were predicted compared to previous work (2003ARA&A..41..429K; 2003AsBio...3..689S; 2015ApJ...806..137R; 2018ApJ...854...19R; 2020ApJ...904...10K; 2023RSOS...1030056J). More recent work has in part supported the results of lower O3 for a given O2 mixing ratio (2022CliPa..18.2421Y; ji2024correlated; liu2025evolution). O3 and its associated UV heating is responsible for the temperature inversion that creates the Earth’s stratosphere, and this heating affects the temperatures in the TTL. Therefore, the O2 fluctuations following the GOE may have modulated the upward transport of H2O through the TTL and consequently, hydrogen escape too.
Using the 3D chemistry-climate model WACCM6, we first explore the affect of changing the O2 abundance on the TTL region, before investigating the influence it may have had on hydrogen escape during the Proterozoic eon. Our aim is to estimate how changes to atmospheric O2 alone can affect the diffusion-limited hydrogen escape rate. In terms of hydrogen escape fluctuations, our focus is on H2O, and we hold the mixing ratios of H2 and CH4 constant at the surface. We note that CO2, CH4 and H2 are likely to have varied during the past 2.4 billion years (2020SciA....6.1420C), as will have the luminosity of the Sun (2012RvGeo..50.2006F). These variables are important because CH4 and H2 are key hydrogen carriers to the upper atmosphere, and additionally CH4 and CO2 provide greenhouse warming which can affect TTL temperatures. The TTL will also have been affected from the reduced total insolation from a younger Sun which would act to cool the troposphere. The purpose of this work is, however, to present the mechanism by which O2 affects hydrogen escape. We calculate such H escape deviations, paving the way for future work on the topic to investigate other physical, chemical, and biological influences.
1 Numerical methods
1.1 WACCM6 model
The simulations we analyse are those from 2022RSOS....911165C, where the Earth System Model WACCM6 (2019JGRD..12412380G) was employed to model different oxygenation levels. WACCM6 is a configuration of the Coupled Earth System Model (CESM), and we use version 2.1.3 (CESM2.1.3222https://www.cesm.ucar.edu/models/cesm2) with minor code alterations in the initial conditions [^3] ExoCESM GitHub: https://github.com/exo-cesm/CESM2.1.3/tree/main/O2_Earth_analogues, which we describe in this section.
The first simulation we perform commences with initial conditions which are intended to be representative of Earth’s pre-industrial (PI) atmosphere circa 1850, alongside contemporary ocean and land settings, and the current orbital setup with a $23.4^{\circ}$ obliquity and 24-hour rotation rate. Deep time paleoclimate studies would often change the continents and we accept that this is a limitation with the WACCM6 model where continents cannot currently be modified. The circulation pattern and climate would have been different in the past due to various factors, which include the younger Sun, different greenhouse gas concentration, and the formation and break up of supercontinents such as Pangea and Columbia/Nuna (Rogers_2002_Columbia; 2012RvGeo..50.2006F; 2017Geo....45..231F; 2020SciA....6.1420C). The model’s spatial resolution for the atmosphere was set at $1.875^{\circ}$ latitude by $2.5^{\circ}$ longitude (yielding a $96\times 144$ horizontal grid). This configuration utilizes 70 atmospheric layers in vertical pressure coordinates extending from the surface (approximately 1000 hPa) to the thermosphere ($4.5\times 10^{-6}$ hPa, or $\sim 150$ km). The land component (CLM5.0) operates on the same $1.9^{\circ}\times 2.5^{\circ}$ horizontal grid as the atmosphere to maintain consistency in surface flux exchanges. The simulation employs an interactive ocean via the POP2 model, which operates on a $1^{\circ}$ displaced-pole grid. This ocean grid features $320\times 384$ points, with enhanced latitudinal resolution near the equator to better resolve tropical dynamics. This same grid is utilized by the Sea-Ice (CICE5) and Wave (WaveWatch3) components. Additionally, freshwater discharge is included in the MOSART river transport model on a $0.5^{\circ}$ global grid.
Both atmospheric and oceanic components of the model are interactive, allowing dynamic responses to environmental stimuli like temperature changes. Our simulations incorporated middle atmosphere chemistry as described by 2020JAMES..1201882E, featuring 98 chemical species and 298 reactions, including O3 chemistry. The atmospheric time step, $\Delta t$, was set at 30 minutes. The concentrations of 75 species were computed using an implicit numerical scheme, while 22 long-lived species were computed using an explicit numerical scheme (sandu1997benchmarking; 2005ama..book.....B). N2 is considered invariant. Absorption of light by CO2 and H2O in the Schumann-Runge (S-R bands) is not included in these simulations, and neither is scattering at wavelengths <200<200 nm. For more details on the inclusion of these processes, we refer the reader to ji2023comparison and ji2024correlated, as well as Section 3.2 where we discuss how it might impact our results.
In total, eight simulations were performed for the Earth system with different mixing ratios of atmospheric O2, where the atmospheric pressure was kept the same by changing the abundance of the background gas, N2. O2 mixing ratios are given in terms of present atmospheric level (PAL), which is 21% by volume. The O2 surface mixing ratios span values between 0.1 – 150% PAL, which are within estimates for the Earth over the last 2.4 billion years (2014Natur.506..307L; 2019MinDe..54..485L; 2020SciA....6.1420C; 2020PreR..34305722S). These temporal estimates do not always agree, so we remain agnostic and perform simulations covering the whole range. The simulation names are PI, 150% PAL, 50% PAL, 10% PAL, 5% PAL, 1% PAL, 0.5% PAL, and 0.1% PAL. CO2, CH4, N2O, and H2, were held fixed at values of 284 ppmv, 0.8 ppmv, 273 ppbv, and 0.5 ppmv, respectively. The mixing ratios or fluxes of gases specified at the surface were kept constant as O2 was altered. This assumption is likely not realistic for the entirety of the last 2.4 Gyr, but it allows us to isolate the role O2 changes make to the chemistry and dynamics of the atmosphere. Water vapor feedback, in terms of evaporation and rainout, is included in WACCM6, but hydrogen escape and water loss are not explicitly simulated. Instead, we estimate these values once quasi-steady state has been reached in the simulation (see section 1.2).
We average the outputs of the WACCM6 simulations over the last four years of the simulation once the middle atmospheric (defined here as between 12 – 100 km in altitude; 2005ama..book.....B) trend in total hydrogen atoms has halted, and this required between 30 and 90 simulations years for the different scenarios.
1.2 Diffusion-limited hydrogen escape rate
The diffusion-limited hydrogen escape rate (1973JAtS...30.1481H; 2003ARA&A..41..429K), $\Phi_{\textrm{esc}}$, is proportional to the total mixing ratio of hydrogen-bearing species, $f_{\textrm{T}}(\textrm{H})$, at the homopause:
$\Phi_{\textrm{esc}}\propto f_{\textrm{T}}(\textrm{H}),$ (1)
where $f_{\textrm{T}}(\textrm{H})$, can be written as
$f_{\textrm{T}}(\textrm{H})=f(\textrm{H})+2f(\textrm{H}_{2})+2f(\textrm{H}_{2}\textrm{O})+4f(\textrm{CH}_{4})$..., (2)
and so on, accounting for all hydrogen-bearing species. We focus on evaluating the abundances of H, H2, H2O, and CH4. In modern Earth’s atmosphere, at the tropopause, the H2 and CH4 mixing ratios are $\approx 0.55$ ppmv (catling2017atmospheric) and $\approx 1.9$ ppmv (Saunois_Methane_2025), respectively. There, the H2O mixing ratio varies between $\approx 2.5-4.5$ ppmv, with a mean stratospheric entry value of $\sim 3.5-4$ ppmv (2009RvGeo..47.1004F; 2019JGRD..12412380G). This puts the stratospheric entry value of total hydrogen at $\sim 17$ ppmv, although with CH4 at a lower concentration during the pre-industrial period, the entry value would be closer to $\sim 12$ ppmv. In our simulations, the lower boundary mixing ratios are instead mixing ratios of 0.5 ppmv and $0.8$ ppmv for H2 and CH4, respectively, corresponding to pre-industrial values. Consequently, H2O is the primary carrier of hydrogen atoms to the middle atmosphere. Other H-bearing species such as OH, HO2, and H2O2 are included in the WACCM6 model but are not sufficiently abundant to contribute significantly to the overall hydrogen budget.
The calculation of $\Phi_{\textrm{esc}}$ at the homopause can be made by considering the binary diffusion coefficients which depend on the temperature of the atmosphere and the chemical species. $\Phi_{\textrm{esc}}$, in units of atoms cm-2 s-1, is given by
$\Phi_{\textrm{esc}}=\sum_{i=1}^{n}\frac{N_{i}f(i)b_{i}}{H_{h}},$ (3)
where $H_{h}$ is the atmospheric scale height at the homopause, $N_{i}$ is the number of hydrogen atoms in each chemical species, $i$, and $b_{\textrm{i}}$ is the binary diffusion parameter for each constituent. The binary diffusion coefficient depends on the temperature of the atmosphere at the homopause ($T_{h}$), the masses of the species, and their collision cross-sections. For each species considered here, the binary diffusion coefficients are $b_{\textrm{H}}=6.5\times 10^{17}T_{h}^{0.7}$, $b_{\textrm{H}_{2}}=2.67\times 10^{17}T_{h}^{0.75}$, $b_{\textrm{H}_{2}\textrm{O}}=0.137\times 10^{17}T_{h}^{1.072}$, $b_{\textrm{CH}_{4}}=0.734\times 10^{17}T_{h}^{0.750}$. These binary diffusion parameters for H, H2, and H2O are sourced from 1973JAtS...30.1481H and CH4 is sourced from banks1973aeronomy. The calculations are made for the whole atmosphere (latitude and longitude at the homopause) before calculating a longitudinally-averaged and latitudinally-weighted mean.
From Earth’s exosphere, hydrogen escapes faster than it is delivered, such that throughout this work, we will assume that the loss rate of hydrogen atoms is diffusion-limited (ignoring hydrodynamic escape and impacts). Note that it is possible for hydrogen to escape slower than this upper limit, provided the rate of loss from the exosphere is the limiting factor (e.g., energy-limited escape, photon-limited escape; 2005Sci...308.1014T; 2016A&A...585L...2S; 2016ApJ...816...34O; 2017ApJ...845..130L).
2 Results
Figure 1: The zonal mean of atmospheric temperature (in K) is shown for simulations with O2 mixing ratios between 150% PAL and 0.1% PAL, where O2 decreases in abundance from the top left panel to the bottom right panel. The colour bar shows hotter temperatures in red and colder temperatures in blue. The TTL region is $\pm 23.4^{\circ}$ from the equator and between 70–150 hPa in pressure.
As part of the hydrological cycle, liquid water evaporates from the surface and warm, moist air rises due to buoyancy. Hence, rivers, lakes, and predominantly the ocean, act as a sources of hydrogen atoms to the atmosphere. We will present the results in terms of how the total hydrogen mixing ratio starts at the surface and how it changes with decreasing atmospheric pressure (increasing altitude). This includes tracking the proportion of hydrogen atoms bound in different species, such as H2O and CH4. We will show how the O2 mixing ratio affects the total hydrogen mixing ratio through the atmosphere and ultimately the amount of hydrogen atoms diffusing up to the exosphere. In the following subsection, we consider how O2 affects clouds, condensation, and the cold trap at the top of the troposphere. These processes act to keep hydrogen atoms in the troposphere that would otherwise mix into the stratosphere and upwards to eventually escape to space.
2.1 Water vapour through the cold trap
Figure 2: Left: The mean tropical (defined at latitudes $\pm 24^{\circ}$ from the equator) O3 mixing ratio between 70 – 150 hPa is plotted on the left vertical axis in teal circles against the atmospheric mixing ratio of O2 at the surface in terms in terms of the present atmospheric level (PAL), which is 21% by volume. The total global ozone (O3) column density in Dobson Units (DU, where 1 DU $2.69\times 10^{20}$ molecules m-2) is also plotted in purple squares on the right vertical axis against the atmospheric mixing ratio of O2. Right: The O3 number density is shown for all eight simulations between 10 – 200 hPa on a linear scale (note the number density does not go to zero at pressures greater than 200 hPa, see 2022RSOS....911165C).
In every simulation, the global mean mixing ratio of each species at the surface for H2O, H, CH4, and H2, is $\sim 10,000$ ppmv, $10^{-8}$ ppmv, $\approx 0.8$ ppmv, and $\approx 0.5$ ppmv respectively, with the latter two molecules specified with fixed mixing ratios at the surface. As temperature and pressure decline with altitude in the troposphere, the atmosphere cannot hold as much H2O vapor. This is because the drop in pressure causes the air to cool (adiabatic cooling), and the saturation vapor pressure of water, which determines the air’s capacity for water vapor, is strongly dependent on temperature. Fig. 1 presents the zonal mean (averaged over longitude) of the atmospheric temperature between the surface and 10 hPa ($\sim 30$ km in altitude) for all the WACCM6 simulations. The temperature structure of the atmosphere is set by the strength and spectral shape of the incoming radiation, the moist adiabat in the troposphere, atmospheric composition (providing heating and cooling in different layers), clouds and ice, and atmospheric dynamics. Between 150% PAL and 50% PAL, the temperature structure at pressures greater than 10 hPa (or altitudes below 10 hPa) is similar. At 10% and 5% PAL, the temperature increases in the middle atmosphere, and this is noticeable in the tropics between 200 – 30 hPa and at northern ($45-90^{\circ}$) latitudes between 200 – 100 hPa. At $\leq 1$% PAL of O2, the stratosphere begins to become significantly cooler, with the 10 hPa pressure level $\sim 60$ K cooler in the 0.1% PAL simulation compared to the 150% PAL simulation. A globally averaged stratospheric temperature inversion of $\approx 2$ K and $\approx 1$ K exists in the 1% PAL and 0.1% PAL simulations, respectively, which is significantly smaller than the $\approx 60$ K temperature inversion in the PI atmosphere.
Fig. 2 shows the mixing ratio of atmospheric O2 plotted against the O3 column density, as well as the mean O3 mixing ratio between 70 and 150 hPa in the tropical region, accounting for the expanse of the TTL. The figure also shows the O3 number density in the tropics between 10 and 200 hPa. Whilst the global O3 column density increases with increasing O2 mixing ratio, this is not the case for the O3 mixing ratio around the TTL region: there is instead a peak at $\approx 5-10$% PAL. The O3 number density also mirrors this relationship. When O2 is decreased, O3 accumulates lower down in the atmosphere because the wavelengths that photolyse O2 to produce “odd oxygen” (O + O3) are able to penetrate deeper. Consequently, for O2 concentrations of $\leq 10$% PAL, O3 maximizes in this region near the TTL. The TTL temperatures are a balance between shortwave heating and the emission and absorption of longwave radiation. Whilst this causes higher temperatures than the PI atmosphere for 5 – 10% PAL, simulations with $\leq 1$% PAL have resultant lower temperatures. This effect is a local TTL phenomenon; globally, the mean tropopause temperatures decrease with decreasing O2.
These results are in line with previous O3 depletion experiments (albeit of a much smaller magnitude) which found that reduced O3 concentrations cooled the TTL and led to less stratospheric water vapour (Forster_2007_O3; Xie_2008_O3_tape; Lu_2021_O3; Zhou_2022_Ozone). The cooling was attributed to radiative processes. Indeed, Forster_2007_O3 showed that less O3 in the tropical lower stratosphere caused cooling that extended into the upper tropical troposphere due to reduced downward longwave radiation. Similarly, here we find that lower middle stratospheric densities of O3 reduces the propagation of longwave radiation downwards. This mechanism is the cause of the hotter temperatures in the higher oxygenated scenarios in Fig. 1 compared to scenarios with $\leq 1$% PAL. This is also why there is an asymmetry in temperatures because of hemispherical asymmetries in the O3 distribution (there is more O3 in the northern hemisphere) as a result of the Brewer-Dobson circulation (1926RSPSA.110..660D; 1949QJRMS..75..351B; 2014RvGeo..52..157B).
In Section The Oxygen Valve on Hydrogen Escape Since the Great Oxidation Event, we described how H2O moves through the TTL to get into the stratosphere on modern Earth. The cold trap mechanism is indirectly sensitive to O2 concentrations resulting from the aforementioned difference in O3 and hence UV heating. Due to the difference in time-averaged zonal mean TTL temperatures between the PI and 0.1% PAL simulations of up to $\approx 17$ K, more H2O is frozen out in the form of water clouds, ice, and ice clouds (e.g., cirrus clouds; 2012JGRD..117.4211W) before crossing $\approx 100$ hPa in the 0.1% PAL case compared to the standard PI case. Stratospheric O3 and H2O are greenhouse gases, and with less H2O reaching the stratosphere as a consequence of reduced O3, this decreases tropospheric heating (2023CliPa..19.1201D), causing a positive feedback effect, albeit a small one: there is some cooling at the surface, but this is limited to <3<3 K in terms of global averaged surface temperatures.
The fractional coverage of high clouds (high clouds are defined as clouds at pressures <400<400 hPa and are therefore relevant to the TTL) versus latitude is shown in Fig. 3. The coverage of high clouds for the PI, 1% PAL, and 0.1% PAL is also shown in the same figure. The O2 mixing ratio affects high cloud coverage across the entire planet, but especially in the tropics. There is a $\approx 30$% increase in the peak high cloud fraction at the equator ($0^{\circ}$ latitude) from the 150% PAL simulation to the 0.1% PAL simulation, although the peak high cloud fraction for the simulations between the O2 concentrations of 150% PAL to 1% PAL at the equator all fall within 6% of each other. Low clouds (surface – 700 hPa) and medium clouds (700 – 400 hPa) exhibit very little deviation between the simulations.
[…]
Figure 3: The top left, top right, and bottom left panels show the high cloud fraction as a function of longitude and latitude for the PI, 1% PAL, and 0.1% PAL scenarios, respectively. Bottom right: The high cloud fraction (high clouds are defined as clouds at pressures <400<400 hPa) is averaged over longitude and shown as a function of latitude for all the WACCM6 simulations.
Figure 4: The atmospheric ice content (blue-white shading) and the amount of atmospheric H2O (magenta-orange-yellow contours) is shown on the latitude-longitude grid of the Earth. Both are given in terms of ppmv. The PI simulation (left) and the 0.1% PAL simulation are shown (right). Both scenarios are shown for a pressure level of 142 hPa, which corresponds to an altitude of $\approx 14$ km near the base of the TTL.
In the cases we have simulated, the interplay between UV radiation, O2 mixing ratio, and 3D transport, is especially important for predicting the O3 distribution. The amount of O3 and the incoming UV affect temperatures around the tropopause, and ultimately how much H2O can reach the stratosphere. Here we are exploring the impact of O2 levels alone on the TTL temperatures and thus atmospheric escape. Other factors will have influenced TTL temperatures for the past 2.4 billion years, and these will be discussed in Section 3.
2.2 Tropical tape recorder
The tropical “tape recorder” (1995GeoRL..22.1093M; 2007GeoRL..34.9804L; 2017ACP....17.4337G) for H2O is the phenomenon where seasonal temperature fluctuations in the TTL modulate the mixing ratio of H2O that enters the stratosphere over an annual cycle. The temperature of the TTL itself is affected by several processes, including radiative heating and cooling, deep convection, sea surface temperatures, and the Quasi-Biennial Oscillation (2009RvGeo..47.1004F; 2012ACP....1212183P; 2013JGRD..118.9658G; 2020GeoRL..4789533T). Note that the tape recorder effect exists for some other molecules too, including CO (2006GeoRL..3312811S; 2007GeoRL..34.9804L; 2008GeoRL..35.5801P). The Brewer-Dobson circulation (1926RSPSA.110..660D; 1949QJRMS..75..351B; 2014RvGeo..52..157B), over a period of months, transports H2O upward (2017ACP....17.4337G), causing the distinct tape recorder effect (WACCM6 is able to generate this effect - see the PI simulation in Fig. 5 and also 2019JGRD..12412380G). The seasonally-varying temperatures in the TTL imprint a record on the observed mixing ratio of H2O which is altered in the middle atmosphere in the subsequent months.
In the PI simulation, the tape recorder is easily visible over a 4-year time frame with a period of 1 year (see Fig. 5). For the PI case, the peak seasonal differences in the tape recorder effect are of order 1.5 ppmv at $\approx 90$ hPa. Note that this amplitude is much smaller than the tape recorder variation 2023MNRAS.524.1491L found for an Earth-like exoplanet with a higher eccentricity of 0.4.
At low oxygenation states (<1<1% PAL), seasonal differences in temperature and the Brewer-Dobson circulation are still present, but their magnitude is smaller so the tropical tape recorder effect is muted. Whilst the H2O mixing ratios still vary, the tape recorder effect is no longer visible from the scale shown in Fig. 5 at 0.1% PAL because there is no clear periodic signal. Variance-preserving Fourier spectra show that the annual harmonic (1 cycle yr-1) dominates the stratospheric H2O variability in the PI simulation between 100 and 30 hPa. In contrast, the 0.1% PAL simulation exhibits a much weaker annual contribution relative to total variance at these pressure levels, consistent with the absence of a tape-recorder signal. Therefore, the seasonal mechanism which transports more water vapour into the stratosphere during warmer periods in the TTL is no longer effective in the 0.1% PAL case.
Figure 5: The water vapour mixing ratio in ppmv is shown for the PI scenario (left) and the 0.1% PAL scenario (right) between pressures of 120 – 30 hPa. Each panel shows four years in a row of each simulation, displaying the zonally averaged and latitudinally weighted H2O mixing ratio in the tropics between $\pm 24^{\circ}$. Whites show the lowest mixing ratios and darker blues show progressively larger mixing ratios. Note the different scales on each colour bar. The contours show the atmospheric temperature in kelvin, with yellow indicating higher temperatures and magenta indicating lower temperatures.
2.3 Total hydrogen mixing ratio
After H2O passes through the cold trap, the first bottleneck for the total hydrogen mixing ratio of temperate Earth-like atmospheres has been passed. We now assess the quantitative effect of this bottleneck, and the relative contributions of each of the four major hydrogen bearing species in our simulations.
Figure 6: Left: The globally averaged total hydrogen mixing ratio ($f_{\textrm{T}}(\textrm{H})$) at 50 hPa is plotted against the warmest temperature reached within $\pm 24^{\circ}$ (tropical latitudes) from the equator at 88 hPa for each simulation. Linear regression gives a coefficient of determination ($R^{2}$) of 0.994, which is showed by the dotted black line. The different O2 mixing ratios are shown by various markers: 150% PAL (grey cross), PI (black square), 50% PAL (dark green plus), 10% PAL (green downward triangle), 5% PAL (light green star), 1% PAL (dark blue circle), 0.5% PAL (blue diamond), and 0.1% PAL (light blue upward triangle). Right: The H2O mixing ratio is plotted against pressure for the PI (black) and 0.1% PAL (light blue) simulations, in terms of its tropical (unbroken line) and global (dotted line) average. Also shown is the ice saturation mixing ratio which depends on temperature and pressure and is shown for the tropical (dashed) and global (dash-dotted) average. The tropical average is for latitudes $\pm 24^{\circ}$ from the equator.
Fig. 6 displays the warmest atmospheric temperature reached in each simulation at a pressure of 88 hPa ($\approx 18$ km for Earth’s atmosphere)44488 hPa is the pressure level in our simulations closest to the winter climatic tropopause of 86-88 hPa, although note that the summer climatic tropopause is closer to 100 hPa (kim2012tropical; pilch2016tropical) in the tropics (within $\pm 24^{\circ}$ from the equator) against the globally averaged $f_{\textrm{T}}(\textrm{H})$ at the same pressure level. There is a strong positive correlation with a coefficient of determination of $R^{2}=0.994$ between these two variables. This freeze-drying relationship arises because the warmest temperatures in the tropical atmosphere effectively control the maximum amount of water vapor that can be transported upward through to the stratosphere. As a result, there is a weaker correlation present when comparing $f_{\textrm{T}}(\textrm{H})$ with the global mean temperature at the same pressure level, and worse correlation still when comparing to the warmest temperature found anywhere at 88 hPa.
More importantly, Fig. 6 also demonstrates how it is the ice saturation vapor pressure (murphy2005review) in the tropical atmosphere that acts to limit the transport of H2O upwards. The tropical ice saturation mixing ratio curve is within a factor of 1.18 and 1.01 when compared to the globally averaged H2O mixing ratio at 100 hPa ($\approx 16$ km in altitude) for the PI and 0.1% PAL atmospheres, respectively. When comparing the global ice saturation mixing ratio curve, the discrepancy is a factor of 5.8 and 3.9 times, respectively. This means that 1D models that are based on globally averaged temperatures may overestimate the stratospheric H2O abundance for Earth-like oxygenated simulations by up to a factor of 6 (accounting for all simulations here, the discrepancy range is 3.8 – 6.0).
Fig. 6 suggests that the warmest TTL temperatures are the controlling factor in each atmospheric scenario, instead of the atmospheric composition. Yet because composition (i.e., the O2 mixing ratio) is the variable that is altered in each scenario, the atmospheric composition is the controlling factor for the TTL temperatures. Hence, the oxygenation state of the atmosphere is indirectly controlling the upward flow of hydrogen atoms and affecting the diffusion-limited hydrogen escape rate, and this is what we refer to as the ‘oxygen valve’.
Fig. 7 presents the globally averaged $f_{\textrm{T}}(\textrm{H})$ vertical profile throughout the atmosphere, which is calculated from Eq. 2 for all the simulations. The mixing ratio of total hydrogen in the stratosphere (at 1 hPa) is $f_{\textrm{T}}(\textrm{H})=11.7$ ppmv in the PI simulation. Turbulent mixing causes $f_{\textrm{T}}(\textrm{H})$ to remain roughly constant between the lower stratosphere and homopause. For the 10% PAL, 1% PAL, and 0.1% PAL simulations, the stratospheric $f_{\textrm{T}}(\textrm{H})$ values at 1 hPa are approximately 12.4, 6.9, 3.0 ppmv, respectively.
Fig. 7 also shows the total hydrogen contribution from the four species that carry the majority of hydrogen atoms (H, H2, H2O, and CH4). These panels show that for the PI simulation (black lines), H2O is the primary carrier of hydrogen atoms in the lower atmosphere. Above the cold trap in the TTL, the H2O mixing ratio increases until it reaches a maximum mesospheric value ($\approx 5$ ppmv) due to CH4 reacting with OH. At a pressure of approximately 0.01 hPa, H2 then becomes dominant. Photolysis of H2O, CH4, and H2, as well as diffusive separation, cause atomic H to become the primary carrier at pressures less than $10^{-4}$ hPa.
Decreasing levels of O2 act to shift the pressures at which particular species are the principal hydrogen carrier. H2 begins to increase in abundance at lower altitudes as the atmospheric O2 concentration is decreased. For instance, with 1000 times less O2 (0.1% PAL case, light blue lines), H2 is dominant at pressures less than 10 hPa, due to both H2O condensation and photolysis at the higher pressure levels. Again, resulting from diffusive separation, H always dominates towards the top of the model in the lower thermosphere. CH4 is increasingly lost in the troposphere due to enhanced H2O photolysis and OH production, acting to reduce CH4 lifetimes (see the following papers for a more in depth discussion on CH4 lifetimes: 2022RSOS....911165C; ji2023comparison; ji2024correlated). This loss adds to tropospheric H2O which is ultimately halted by the cold trap. If the CH4 mixing ratio at the TTL is greater than half that of H2O, then CH4 is the primary contributor to hydrogen escape. However, CH4 is never the main H carrier in the scenarios we present. As we will discuss later, it is unclear whether this would have been the case for some portions of the Proterozoic.
Figure 7: Top: $f_{\textrm{T}}(H)$, the total hydrogen mixing ratio is given in parts per million by volume (ppmv) and is shown against pressure for the 150% PAL (grey), PI (black), 50% PAL (dark green), 10% PAL (green), 5% PAL (light green), 1% PAL (dark blue), 0.5% PAL (blue), and 0.1% PAL (light blue) simulations. $f_{\textrm{T}}(H)$ is calculated using Eq. (2). Hence, the mixing ratio contribution from H, H2O, H2, and CH4 is multiplied by a factor of 1, 2, 2, and 4, respectively. Bottom left: The total hydrogen mixing ratio contribution from H (dash dotted), H2 (dashed), H2O (unbroken), and CH4 (dotted) is shown for the PI and 10% PAL simulations and bottom right: the 1% PAL and 0.1% PAL simulations.
2.4 Diffusion-limited escape
Now we have assessed $f_{\textrm{T}}(\textrm{H})$ throughout the atmosphere we can predict how the rate of diffusion-limited hydrogen escape varies between the different simulated atmospheres with WACCM6. We present the predicted escape rates for each simulation in Table 1.
Table 1: This table shows the eight simulations where the mixing ratio of O2, given in terms of present atmospheric level (PAL), was varied. The total hydrogen mixing ratio, $f_{\textrm{T}}(\textrm{H})$, in ppmv at the homopause, which is taken to be at a pressure of $10^{-3}$ hPa. The diffusion-limited hydrogen escape rate is given by $\Phi_{\textrm{esc}}$ in mol yr-1 and compared to the predicted escape rate in the PI simulation ($\Phi_{\textrm{esc,PI}}$).
Simulation O2 [PAL] Homopause $f_{\textrm{T}}(\textrm{H})$ [ppmv] $\Phi_{\textrm{esc}}$ [mol yr-1] $\Phi_{\textrm{esc}}$$\Phi_{\textrm{esc,PI}}$
150% PAL 1.500 9.68 $7.71\times 10^{10}$ 1.02
PI 1.000 9.52 $7.54\times 10^{10}$ 1.00
50% PAL 0.500 10.05 $7.79\times 10^{10}$ 1.03
10% PAL 0.100 12.74 $9.04\times 10^{10}$ 1.20
5% PAL 0.050 13.07 $8.95\times 10^{10}$ 1.19
1% PAL 0.010 8.45 $5.11\times 10^{10}$ 0.68
0.5% PAL 0.005 6.92 $3.89\times 10^{10}$ 0.57
0.1% PAL 0.001 4.09 $1.94\times 10^{10}$ 0.26
The $f_{\textrm{T}}(\textrm{H})$ profile shows that for the cases with $\leq$ 1% PAL of O2, the rate of hydrogen escape is lower than that found in the PI case ($\Phi_{\textrm{esc,PI}}$). The rates for the 0.1% PAL, 0.5% PAL, and 1% PAL simulations are: $0.26\ \Phi_{\textrm{esc,PI}}$, $0.57\ \Phi_{\textrm{esc,PI}}$, and $0.68\ \Phi_{\textrm{esc,PI}}$, respectively. We also predict that the escape rate could have been slightly higher when O2 levels were both greater (e.g. 150% PAL) and lesser (e.g. 5 – 10% PAL) than the present day value, with the maximum predicted hydrogen escape rate in the 10% PAL simulation at $1.20\ \Phi_{\textrm{esc,PI}}$. This shows that the TTL heating and predicted escape rate has a non-linear behavior with O2 concentration. Following the Great Oxidation Event, Earth’s O2 levels are thought to have fluctuated between 0.1% PAL and 150% PAL, although the precise timings and magnitude of these fluctuations are debated (2020SciA....6.1420C; 2020PreR..34305722S; 2020AsBio..20..628P; yierpan2020recycled; krause2022extreme; 2023AREPS..51..253M; 2023GSAB..135..753X; 2024NatCo..15.6794F; 2024NatGe..17..667S), with these fluctuations being due to various physical factors, such as carbonate amassing in the crust (2024NatGe..17..458A), and large igneous province volcanism combined with weathering (2024ComEE...5..609L).
3 Discussion
When the Great Oxidation Event occurred approximately 2.4 billion years ago, O2 increased by several orders of magnitude in the atmosphere in a time frame of approximately 200 Myr (2017PNAS..114.1811G). For the first time, these higher levels of O2 resulted in the generation of a significant O3 layer which may have created a stratospheric thermal inversion (kasting1985oxidant; 2003AsBio...3..689S; 2022RSOS....911165C). The O3 layer provided an additional shield for Earth’s surface from ultraviolet radiation. However, the exact thickness of the O3 layer over geological time is difficult to determine. Many uncertainties exist upon the atmospheric pressure and composition throughout the Proterozoic (2014Natur.506..307L; 2020PreR..34305722S; 2020SciA....6.1420C), as well as there being differences in model predictions under similar assumptions (2017ApJS..231...12W; 2022RSOS....911165C; 2022CliPa..18.2421Y; 2023RSOS...1030056J; ji2024correlated). With an oxygenated biosphere, the conditions were set for the evolution of oxygen-dependent animals (2020Gbio...18..260C) and the eventual colonization of land (dunn2013evolution). Additionally, the GOE may have coincided with the end of relatively high levels of hydrogen escape (2013ChGeo.362...26Z; 2019GeCoA.244...56Z). Whether other terrestrial worlds go through a similar path of geological and atmospheric evolution is unknown, although it has been modelled under different conditions and assumptions (2018AsBio..18..856G; 2018haex.bookE.189O; 2021NatGe..14..138O; 2021AGUA....200294K; 2022ApJ...933..115K).
Our work represents an initial step towards a quantitative estimate for hydrogen escape since the Proterozoic began 2.4 billion years ago. Overall, our results suggest that hydrogen escape during the Proterozoic could have been between 3.8 times lower and up to 1.2 times greater than the calculated pre-industrial escape rate when accounting for O2 changes only. The present day loss of hydrogen is insignificant on geological timescales, such that the fluctuations in O2 from the start of the Proterozoic to the present day likely did not in itself cause substantial levels of hydrogen escape. In what follows, we discuss how our simulations can be used as a platform for future work towards the goal of reconstructing Earth’s past atmospheric hydrogen escape.
3.1 The oxygen-ozone valve
When only changing the atmospheric O2 (and N2) mixing ratio, as O2 increases from 0.1% PAL to 150% PAL, the total atmospheric O3 column increases (2022RSOS....911165C). Note that increasing quantities of total O3 does not always lead to an increase in temperatures in the TTL, because the location of the O3 abundance peak in the atmosphere, and the magnitude of the peak, can move with O2 mixing ratio. This effect is illustrated by the greater total hydrogen mixing ratio at the homopause in the 10% and 5% PAL cases when compared to the PI case (e.g., see Fig. 1, Fig. 6, and Fig. 7).
Our work shows that O2 levels during the Proterozoic may have partially controlled the diffusion-limited flux for hydrogen escape. However, this will depend on several other factors, such as the continental distribution and the Earth’s obliquity which can affect atmospheric dynamics and seasonal cycles, respectively. The fainter Sun, the albedo of the Earth through time, and various greenhouse gases (including CH4) which can act to cool the stratosphere but warm the surface (2024PNAS..12119228L), will have also impacted TTL temperatures and hydrogen escape rates. For instance, a cooling of the stratosphere can increase the amount of O3, just like the current increasing atmospheric content of CO2 is aiding a possible, albeit complex, recovery of the O3 layer (2002JGRD..107.4049R; fahey2018scientific; 2022WCD.....3..139I; 2021ACP....2111041M).
There exist estimates of many of these aforementioned variables through Earth’s history, but each of them come with uncertainties and vary throughout the Proterozoic. liu2025evolution suggest that the O3 layer could have been kept at low levels due to protracted atmospheric iodine concentrations throughout the Proterozoic. Atmospheric iodine would destroy O3 through catalytic cycles (WACCM6 does not currently include iodine in its chemical mechanism). The computed WACCM6 O3 columns are lower than the baseline estimates from liu2025evolution, but when liu2025evolution included substantial atmospheric iodine (e.g. $20\times$ modern marine iodine concentrations that was emitted to the atmosphere), then their O3 columns are lower than the WACCM6 results. If we incorporated iodine, then this could lead to a reduced rate of predicted hydrogen escape at a given O2 concentration due to diminished O3 heating.
3.2 Model limitations
Our atmospheric predictions are relevant for both Earth’s past and future climate states. However, this statement comes with several caveats. The moist-adiabatic lapse rate (the rate at which the temperature of a parcel of saturated air decreases with altitude as it rises in the atmosphere under adiabatic conditions), strongly depends on temperature. Thus, the total hydrogen mixing ratio in the troposphere of the Earth is determined by the incoming radiation, albedo, and greenhouse effect. For modern Earth, these parameters are relatively well known. We did not attempt to fully account for all climatic changes since the Proterozoic began; instead, as a first step, we focused on the effect on the escape rate by changing O2 only. The range of O2 concentrations we accounted for are relevant to the past 2.4 billion years and are not relevant to the Archean eon (4 – 2.4 Gyr ago) when molecular oxygen was much lower than present day and Proterozoic (2.4 – 0.541 Gyr ago) concentrations (2020SciA....6.1420C). We remind the reader that WACCM6 is designed for the modern Earth and the current set of simulations do not account for a faster rotation rate in the past (1987PreR...37...95Z; 2016GeoRL..43.5716B; laskar2023did) or different continental coverage (2021NRvEE...2..358M). Additionally, we note that changes to the orbital parameters, rotation rate, or obliquity, may affect the tape recorder signal.
Since 2022RSOS....911165C was published, ji2023comparison showed that absorption of light by H2O and CO2 in the S-R bands, as well as scattering in the S-R bands, are increasingly important at lower atmospheric O2 mixing ratios. The simulations presented in this work do not include either of these treatments. Some WACCM6 simulations (at 10% PAL, 1% PAL, and 0.1% PAL) with H2O and CO2 absorption in the S-R bands were included in ji2023comparison, and we also ran some simulations with fixed surface fluxes (e.g. for CH4). Whilst the absolute mixing ratios do deviate when this absorption is included, the effect of changing O2 on temperature is stronger. In terms of scattering, ji2023comparison showed that it is most important at $\leq 1$% PAL of O2. Inclusion of scattering will decrease the total O3 column and likely result in a cooler TTL region. Again, the conclusion that hydrogen escape can be be modified by O2 concentrations remains unaffected, especially when the prescribed CH4 mixing ratio sets a minimum on the hydrogen escape rate because it is not cold-trapped. Nevertheless, future simulations where WACCM6 includes both absorption and scattering in the S-R bands will provide more comprehensive atmospheric calculations.
3.3 Future work for Earth
The different climate states that might have existed throughout the Proterozoic and Phanerozoic [^5] By climate states, here we refer to the variety of factors that would have affected weather, atmospheric circulation systems, and surface temperature. These include glaciations and hotter periods, as well as the various continental land masses, fainter Sun, and greenhouse gas concentrations. (0.541 Gyr ago – present day) should be simulated in future work. A hotter troposphere may have existed due to increased amounts of CH4 or CO2, or both. In this case, the warmer TTL would have increased hydrogen escape rates by elevating the amount of H2O reaching the stratosphere. On the other hand, colder temperatures during Earth’s glacial periods would cool the tropopause (2019JGRD..12411819G) and reduce the stratospheric H2O concentration.
Even with a much colder TTL, CH4 bypases the control of the oxygen valve or a colder troposphere because CH4 does not condense in Earth’s atmosphere like H2O. Thus, the Proterozoic and Phanerozoic hydrogen escape rate could have instead been controlled by the CH4 abundance rather than the O2 abundance, assuming that its mixing ratio reaching the TTL was greater than half the mixing ratio of H2O in the TTL - see Eq. 2. The current problem with evaluating this effect is that estimates for Proterozoic CH4 concentrations vary by several orders of magnitude (2003Geo....31...87P; 2016E&PSL.434...42D; 2018Geo....46..139Z; 2018haex.bookE.189O; 2019E&PSL.522...48L; 2020SciA....6.1420C; cadeau2020carbon) and may be greater or lower than the present day mixing ratio. The thinner simulated O3 columns found from 3D modeling (2022RSOS....911165C; 2022CliPa..18.2421Y) and updated 1D calculations (ji2024correlated), as well as the inclusion of iodine chemistry (liu2025evolution), suggest that more UV radiation would have penetrated deeper into the atmosphere for much of the Proterozoic and potentially resulted in lower CH4 lifetimes at O2 concentrations of $\leq 1$% PAL. In this situation, greater surface-to-atmosphere fluxes of CH4 would have been required to sustain current $\sim 1$ ppmv mixing ratios (such fluxes may not be plausible; daines2016effect; Olson2016LimitedGreenhouse; 2019E&PSL.522...48L), or the higher estimates that some studies suggest in the region of $\sim 10$$100$ ppmv (2003Geo....31...87P; 2017Geo....45..231F; zhao2018terrestrial; fakhraee2019proterozoic). Closer in time to the present day, CH4 mixing ratios in the Phanerozoic (0.51 Gyr ago - present day) have been simulated to be between 0.1 – 12 ppmv (beerling2009methane).
Clearly, there is much to still learn about why the different models predict specific atmospheric abundances for molecules such as O3, which has significant horizontal and vertical variation (2022RSOS....911165C; 2022CliPa..18.2421Y), and CH4 (kasting2025evolution). The entire conceivable parameter space has certainly not been fully explored; such endeavors utilising 3D models are welcomed but will likely prove computationally expensive. Attempts to define the CH4 concentrations in the various geological eons are an important part of the history of hydrogen escape, but for now, the mixing ratio of CH4 through time remains poorly constrained. Ascertaining its production rate in the Proterozoic will help to establish its concentrations in that eon (kasting2025atmospheric).
Eventually, as the Sun’s luminosity increases, the Earth will heat up, CO2 will be sequestered out of the atmosphere through carbonate-silicate weathering, and it is predicted that the Earth will lose all of its water through the moist greenhouse effect approximately 2 Gyr in the future (1984Icar...57..335K; 1988Icar...74..472K; 2014GeoRL..41..167W; 2021NatGe..14..138O), although the estimate of the time varies. But before that event occurs, previous studies have predicted that the lack of CO2 will have a detrimental effect on the biosphere because plants are unable to survive at very low CO2 concentrations (1982Natur.296..561L; 1992Natur.360..721C).
Several calculations of the lifespan of the Earth’s biosphere have been attempted (1982Natur.296..561L; 1992Natur.360..721C; 2000TellB..52...94F; von2003biogenic; 2018AsBio..18..469R; 2021NatGe..14..138O; 2023IJAsB..22..272M; Graham_2024_Extension). For instance, 2021NatGe..14..138O calculated that in $1.08\pm 0.14$ billion years ($1\ \sigma$), Earth’s atmospheric O2 will drop to 1% PAL. More recently, Graham_2024_Extension suggested that Earth’s biosphere may last for up to 1.6 – 1.86 Gyr. Most of these estimates for biospheric termination are before the Earth is predicted to lose its water and become, by the most basic definition, uninhabitable. As O2 levels drop following the extinction of photosynthesizing life (2021NatGe..14..138O), O3 will decrease, and the upper troposphere and TTL may be colder than under otherwise equivalent conditions, diminishing the amount of hydrogen available in the upper atmosphere which can escape to space and cause irreversible water loss (the future rates of hydrogen escape would likely have a negligible effect on oxidizing the Earth to reverse the loss of O2 from ceasing photosynthesis). Hence, our results suggest that the ocean loss timescale could be extended. Thus, investigations aiming to assess the lifetime of Earth’s future habitability should incorporate a 3D chemistry-climate model that includes comprehensive atmospheric oxygen chemistry and that can adequately represent the 3D movement of air parcels through the TTL. Additionally, it is possible that our WACCM6 simulations could be used as input to a model which has a higher top (lower minimum pressure), such as the WACCM-X model (2018JAMES..10..381L), or even used as an input into a computational model that explicitly calculates atmospheric escape (e.g., 2018PNAS..115..260D; 2023MNRAS.523..286S).
3.4 Future work for exoplanets
Other planets outside of our solar system (exoplanets) may be habitable and evolve life. Whether they can keep liquid water on their surface is crucial for the persistence of habitability as it is currently defined.
The H2O tropical tape recorder is a stratospheric signal of seasonality, impressing the varying TTL temperatures on the stratospheric H2O mixing ratio. Prior work showed how a 15% reduction in global O3 (Xie_2008_O3_tape) could cool the TTL and raise it in height, ultimately weakening the tape recorder signal. Here we find this effect exacerbated further as O2 and O3 reduce by orders of magnitude. Transmission spectra of exoplanet atmospheres probe the stratosphere, and whilst beyond the detection limit of today’s telescopes, such chemical variability may be detectable with future technology or where exoplanets exhibit higher levels of variation that Earth (e.g., for exoplanets with eccentricity $\geq 0.4$ 2023MNRAS.524.1491L) or different types of stratospheric variability (Cohen_2022_LASO). Seasonality of biosignatures and other gases (Olson_2018_season), combined with the H2O tape recorder signal, may aid in determining whether a planet is inhabited. Should there be no tape recorder signal due to a tropopause that does not significantly vary with seasons, then this variability will be more difficult to observe, and H2O will be harder to constrain due to a dry middle-upper atmosphere.
Simulations with other 3D models such as the LMD-g (2022CliPa..18.2421Y) or the Unified Model (2022MNRAS.517.2383B), or those at higher resolutions (e.g., LFRic-Atmosphere; 2023GMD....16.5601S), may yield different quantitative results due to varying treatments of convection and water vapour microphysics, as well as alternative radiative transfer schemes. The stratospheric thermal inversion in Earth’s atmosphere is due to O3 which provides sufficient atmospheric heating when O2 is present in quantities $\gtrsim 0.1$% PAL (2017ApJS..231...12W; 2022RSOS....911165C). But a terrestrial exoplanet may not require O2 to modulate the thermal structure: other atmospheric compositions can result in thermal inversions and may similarly act to enhance or reduce water loss. For example, Titan has a thermal inversion due to the prescence of hazes which absorb shortwave radiation (1973Icar...20..437D; 2017aeil.book.....C), and CN on hot super-Earths can also cause a thermal inversion (2021MNRAS.500.2197Z). Additionally, the heating rates will change depending on the distribution of stellar radiation over wavelengths: O3 radiative heating significantly affects planets around F and G stars but less so around M and K stars (2015P&SS..111...62G; 2022A&A...665A.156K; 2024MNRAS.531.1471D). O2 can also build up on exoplanets without the presence of life under specific conditions due to large amounts of CO2 (2007A&A...472..665S) or H2O photolysis (2014ApJ...785L..20W).
There are various planetary system parameters which affect the climate state of exoplanets. Crucial to the atmospheric dynamics is the planetary rotation rate and whether it is tidally locked (2014RSPTA.37230084F; 2016A&A...596A.112T; 2017A&A...601A.120B; 2019AsBio..19...99D; 2025PSJ.....6....5B). The circulation affects the atmospheric temperature structure and how tracers such as H2O are advected. Other salient factors include the strength and spectral shape of the incoming radiation (2020A&A...639A..99E), the continental distribution or lack thereof (2018ApJ...854..171L; 2021ApJ...910L...8Z; 2022MNRAS.513.2761M), obliquity (2019ApJ...877L...6K), eccentricity (2023MNRAS.524.1491L), flares (2021NatAs...5..298C; 2023MNRAS.518.2472R), simulation resolution (2020ApJ...894...84S; 2021ApJ...913..101L; 2022ApJ...934..149S; 2022ApJ...940...87K; 2023NatAs...7.1070Y; Sergeev_2024_Impact; 2024ApJ...965....5G), as well as the planetary Bond albedo and surface composition (2011ApJ...729...54C; 2019ApJ...884...75D; 2020MNRAS.495....1M). Ocean salinity can also be important in modulating the climate of rocky planets (2022GeoRL..4995748O).
By connecting exoplanet simulations to a model that includes the ionosphere (e.g., WACCM-X), one could attempt to answer interesting questions regarding the upper atmospheres of terrestrial exoplanets. For instance, how do the ionospheres of various exoplanets compare to Earth? Just like flares disturb the atmospheres of terrestrial planets (2010AsBio..10..751S; 2019AsBio..19...64T; 2021NatAs...5..298C; 2023MNRAS.518.2472R), do flares (especially for active M dwarf stars) impact their ionospheres differently from those that strike the Earth (2009Ge&Ae..49..983L; 2018SpWea..16.1363S; 2021SoPh..296..157H)? And could those differences manifest in a way which is detectable with future instruments (2018NatAs...2..287M; 2019A&G....60a1.25M)?
The maximum difference we found in the diffusion-limited hydrogen escape between alternative oxygenation states when all other initial conditions and boundary conditions are kept constant is a factor of 4.7. Whilst this change is not significant for Earth under the assumptions made in this work, it potentially could be significant when the ocean loss timescale for a habitable terrestrial exoplanet is similar to a time that a planet could theoretically spend in the habitable zone (i.e. a planet with a 10 Gyr ocean loss timescale with a different oxygenation state could instead have an ocean loss timescale of $\approx 2$ Gyr).
4 Conclusions
Since the dawn of the Proterozoic (2.4 Gyr ago), several properties of the Earth have changed. The continental configuration has shifted several times, the Sun’s luminosity has increased, and the concentration of atmospheric gases such as O2, CO2, and CH4 have fluctuated.
This work used WACCM6 simulations of the Earth to explore the effect that the changing O2 mixing ratio had on the total hydrogen mixing ratio at the homopause in order to predict the possible diffusion-limited hydrogen escape rate through geological time. A 3D model with chemistry and a global representation of water vapour microphysics is critical to simulating the transport of H2O into the stratosphere and thus estimating diffusion-limited hydrogen escape when other atmospheric constituents (e.g., CH4 and H2, which we held constant at the surface) are not the dominant hydrogen bearing species. In our simulated scenarios we showed that 1D models may predict a stratospheric water vapour discrepancy of between a factor of 3.8 – 6.0 depending on the O2 mixing ratio when compared to 3D calculations.
We found that the atmospheric O2 mixing ratio acts as a nonlinear valve on the total hydrogen mixing ratio in the stratosphere. The O3 mixing ratio (which is affected by the abundance of O2) around the tropical tropopause layer (TTL) determines the amount of UV heating that warms the TTL and controls the H2O mixing ratio entering the stratosphere. The obliquity of Earth induces seasonal cycles, which affect phenomena such as the Brewer-Dobson circulation and the tropical tape recorder. We showed that the tropical tape recorder disappeared at 0.1% PAL, which suggests that the present-day seasonal transport of H2O molecules to the stratosphere may not have taken place at low O2 mixing ratios.
Between 0.1% PAL and 150% PAL of O2, the upward diffusion of H2O is modulated in a non-linear manner, with 5% and 10% PAL having the maximum predicted diffusion-limited escape rate of hydrogen due to O3 concentrations causing localized heating. The predicted escape rate changes by up to a factor of 4.7 between all of the simulations, meaning that whilst the oxygen valve affects diffusion-limited hydrogen escape, in the absence of other factors, it would not have resulted in relatively high levels of hydrogen escape and thus water loss since the GOE.
In order to dissect Earth’s atmospheric history, future work should investigate multiple other physical, chemical, and biological factors that may affect the hydrogen escape rate whilst also accounting for the oxygen valve we presented.
[…]
WACCM6 is a publicly available code. The specific release used in this paper was CESM2.1.3, which can be downloaded from the following CESM downloads page. Our code modifications are detailed here in the ExoCESM GitHub.
[…]
The time-averaged data is currently available in the Dryad data repository.
[…]
GJC set up and ran the simulations, performed the analysis, wrote the analysis and plotting code, and wrote the manuscript. DRM and CW supervised the project and provided comments on the manuscript. FSM and MB provided comments on the manuscript.
[…]
The authors declare no competing interests
Acknowledgements.
We would like to thank the reviewers for their comments which helped to improve the manuscript. G.J.C. thanks the Science and Technology Facilities Council for financial support during the PhD when the simulations were conducted (grant number ST/T506230/1). G.J.C. thanks James Rogers and Oli Shorttle for helpful discussions regarding Earth’s history and hydrogen escape. F. Sainsbury-Martinez and C. Walsh would like to thank UK Research and Innovation for support under grant number MR/Z00019X/1 Additionally, C. Walsh would like to thank the University of Leeds and the Science and Technology Facilities Council for financial support (ST/X001016/1). M.B. appreciates support from a CSH Postdoctoral Fellowship. This work was undertaken on ARC4, part of the High Performance Computing facilities at the University of Leeds, UK. The CESM project is supported primarily by the U.S. National Science Foundation.